A cyclist bikes x distance at 10 miles per hour and returns over the same path at 8 miles per hour. What is the cyclist’s average rate for the round trip in miles per hour?
I got 9.0 but the answer is 8.9

Answer by jainenderkapoor(61)

It is a question of average Speed. It has to be solved in three parts.
1. Going from A to B --- 2. Going from B to A --- 3. Average Speed
1. Going from A to B
Distance = d1 = x miles
Speed = s1 = 10 miles per hour
Time = t1 = Distance/ Speed = x/10
2. Going from B to A
Distance = d2 = x miles
Speed = s2 = 8 miles per hour
Time = t2 = Distance/ Speed = x/8
3. Average Speed = Total Distance/Total Time
Total Distance = x + x = 2x
Total Time = x/10 + x/8 = x (1/10 + 1/8) = x ( 4 + 5)/40 = 9x/40
Speed = 2x/(9x/40) = (2x * 40)/9x = 80x/9x = 80/9 = 8.9
HOPE THIS IS CLEAR TO YOU.
IN CASE YOU NEED ANY HELP, I AM AVAILABLE FOR ONLINE TUTORING. YOU CAN CONTACT ME AT kapoorjai1@rediffmail.com